Evaluateſ (1+ - x2) dx using Riemann Sums. {c= cn (CER), i = mat!), Ž= mort1y2nt.). Š P = [porto} i=1
Answer #1


Ane - $ (1+2x-x²) dx using Riemann sume Cam CCER), E NCHED 12 neht (@nu 2013- [hantu} Ex= b-a - lim & f(x *) DX li* = at (DX)